OpenStudy (anonymous):

Kamila plans to build a concrete block wall behind her house. The wall will be 12 feet long, 6 feet high, and 8 inches thick. Each concrete block measures 16 inches long by 8 inches wide by 8 inches deep. How many blocks will Kamila need to build the wall? Each concrete block has two holes through the middle of the block. The holes are 4 inches wide by 4.5 inches long. What is the volume of one concrete block? Each concrete block weighs 36.5 pounds. What is the total weight of the concrete wall? If each block is solid concrete, how much would one block weigh?

OpenStudy (anonymous):

http://prntscr.com/ascpe6 Here is the image. I really need help on this one.

OpenStudy (aaronq):

Since the thickness of the concrete block is equal to the depth required we can ignore the depth (because only 1 block is needed to achieve it) and focus on the surface area. we know that each block is 16 inches by 8 inches in surface area, so all we need to do is find the area of the wall and divide it by the area of the block. First convert to the same units though, 12 ft = 12*12 =144 inches 6 ft = 12*6= 72 inches Area of wall = 144 in * 72 in Area of block = 16 in* 8 in \(\sf \# ~of ~blocks ~needed=\dfrac{Area ~of ~wall}{Area ~of ~block}\)

OpenStudy (anonymous):

WOW Thank you !! I was making it so much harder than it needed to be!

OpenStudy (aaronq):

no problem! i hope you can tackle the rest on your own

OpenStudy (anonymous):

I think I can , If not I will be back! :)

OpenStudy (aaronq):

okay, sounds good :)